OpenStudy (anonymous):

Solve for B. x+2^2+B^2=1^2

OpenStudy (anonymous):

@Nnesha

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

it is clear that \(1^1=1\)?

OpenStudy (anonymous):

yep

OpenStudy (misty1212):

i suppose i meant \(1^2=1\) but whatever it is still 1

OpenStudy (anonymous):

yerp

OpenStudy (misty1212):

kinda hard to read did you mean \[(x+2)^2+B^2=1\]?

OpenStudy (anonymous):

yeah sorry lol

OpenStudy (misty1212):

and you want \(B\) in terms of \(x\) or \(x\) in terms of \(B\)?

OpenStudy (misty1212):

oh nvm it says "solve for B

OpenStudy (misty1212):

\[B=\sqrt{1-(x+2)^2}\] is the simplest answer

OpenStudy (anonymous):

All right thanks thats what i got.

OpenStudy (misty1212):

actually that is not quite correct, should have \(\pm\) before the radical

OpenStudy (anonymous):

the back of my book says. \[\sqrt{-x-4x-3}\]

OpenStudy (anonymous):

Thats what it says the answer is

OpenStudy (anonymous):

@misty1212